OpenStudy (anonymous):

Use the equation of the tangent line at x=2 to approximate f(2.03) for f(x)= ln(cuberoot(x+1)). Your work must include the equation of the tangent line using exact values. Give 3 decimal places in your final answer.

OpenStudy (cwrw238):

can you differentiate f(x) = ln (x+1)^(1/3)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/3x+3

OpenStudy (cwrw238):

right - so that's the slope of the tangent to the curve so can you find the equation of the tangent at x = 2?

OpenStudy (anonymous):

oh ok I see what to do now, I just didn't know what I needed to do. So the equation of the tangent is y=x/9+7.288/9

OpenStudy (anonymous):

then I assume I would just plug in x?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

brilliant, thanks

OpenStudy (anonymous):

I also have another question about inverses. you don't have to answer it for me, I just don't know where to start.- Full question- If h-1 is the inverse function of h, with the values of h and h '(x) given in the table below, find the slope of the line tangent to the graph of y = h-1(x) at x = 3. Table- x h(x) h'(x) 6 3 7 8 4 6

OpenStudy (cwrw238):

sorry gotta go right now

OpenStudy (anonymous):

ok thanks for your help!